# A systematic review of metaheuristics-based and machine learning-driven intrusion detection systems in IoT

### v3

@@ 3 Related Work
The systematic literature review (SLR) in [hu2024deep] covers only a few population-based metaheuristics for intrusion detection in the IoT environment. Additionally, physical law-based, human-inspired, and hybrid optimization techniques are not discussed. Moreover, the authors do not concentrate solely on IoT. Rather, wireless, public networks, computer networks, Hadoop and MapReduce, and edge networks are also explored significantly. Furthermore, though they analyze the datasets used in developing the IDSs, no correlation is discovered among the metaheuristics, mostly used datasets and ML methods. Saadouni et al. [saadouni2024intrusion] presents an SLR for IoT-IDS based on bio-inspired and ML-driven techniques. The authors vividly discuss the integration of ML methods with optimization algorithms. However, one of the most notable limitations of this work is the investigation of only 25 papers, whereas there are several well-established metaheuristics-assisted IDSs dedicated to IoT. Importantly, although the authors claim to study only IoT-based papers, we find that most of the articles are not focused on the IoT environment. In [reddy2024systematic], the authors study population-based optimizations, specifically swarm intelligence devised for detecting intrusions in IoT. Besides, they analyze the datasets used and the performances of the existing systems. The main drawback of the SLR is the lack of covering all categories of metaheuristics-driven IDSs. Moreover, only basic ML algorithms are discussed in the SLR; whereas many crucial deep learning-based IDSs are sorted out in our study. Sharma et al. [sharma2024multi] aims to explore IoT-IDSs that rely on only multi-objective metaheuristics algorithms. Apart from this, they analyze different machine-learning models and popular datasets. However, the relation between these diverse techniques is still missing. Moreover, the study is not systematic. Heidari et al. [heidari2023internet] introduces an SLR containing rigid comparison and exploration of different IDSs in the IoT environment. The first and foremost limitation is the missing metaheuristics and ML-based systems, which we aim to cover in our study. Verma et al. [verma2020machine] propose a survey on ML-driven IDSs for IoT applications. Regrettably, the study does not include any optimization algorithms, rather it intends to analyze the machine learning classifiers commonly utilized in intrusion detection. Hajiheida et al. [hajiheidari2019intrusion] also do not study the metaheuristics algorithms that are extensively employed in the IoT-IDS. Importantly, rather than focusing on the metaheuristics or ML-based systems, they categorize and discuss different systems, such as SIDS, AIDS, SpIDS, and HyIDS.
@@
The existing reviews either explore nature-inspired or ML-based IoT-IDS systems. Though there is only one survey [sharma2024multi] that discusses the integration of metaheuristics and machine learning techniques for the IoT environment, the coverage and analysis are too inadequate regarding the volume, significance, and diversity of the related IDSs in literature. Most importantly, no reviews analyze the correlation between the optimization algorithms and machine learning models while experimenting on a specific dataset. Apart from these, the existing works do not categorize the applications of different metaheuristics regarding intrusion detection in the IoT environment.
@@
To address all these issues, a systematic literature review is presented, which extensively explores the existing metaheuristics and ML-integrated IDSs, specific to the IoT environment. Additionally, we analyze these systems based on different applications of metaheuristics algorithms, such as optimal feature selection, parameter tuning, hyperparameter tuning, etc. Moreover, the discovery of the connections among these algorithms, their outstanding performances, and the used datasets significantly distinguish this review from others. Furthermore, a new large-scale visualized taxonomy is demonstrated to provide researchers with an overview of the existing IoT-IDSs. Table 1 provides an in-depth comparison with the state-of-the-art reviews, highlighting the contributions of our work over others.
@@
| [hu2024deep] | 2024 | ✓ | 145 | ✕ | ✕ | ✓ | ✕ | ✕ | ✕ | ✕ | ✕ | ✕ | ✓ | ✕ | ✕ |
| [saadouni2024intrusion] | 2024 | ✓ | 25 | ✕ | ✓ | ✓ | ✕ | ✕ | ✓ | ✓ | ✕ | ✕ | ✓ | ✕ | ✕ |
| [reddy2024systematic] | 2024 | ✓ | 101 | ✕ | ✕ | ✓ | ✕ | ✕ | ✓ | ✕ | ✕ | ✓ | ✓ | ✕ | ✕ |
| [sharma2024multi] | 2024 | ✕ | 37 | ✓ | ✓ | ✓ | ✕ | ✓ | ✓ | ✓ | ✕ | ✕ | ✕ | ✕ | ✕ |
| [heidari2023internet] | 2023 | ✓ | 24 | ✓ | ✕ | ✕ | ✕ | ✕ | ✕ | ✕ | ✕ | ✕ | ✓ | ✕ | ✕ |
| [verma2020machine] | 2020 | ✕ | 25* | ✓ | ✕ | ✕ | ✕ | ✕ | ✓ | ✕ | ✕ | ✕ | ✓ | ✕ | ✕ |
| [hajiheidari2019intrusion] | 2019 | ✓ | 43 | ✓ | ✕ | ✕ | ✕ | ✕ | ✓ | ✓ | ✕ | ✕ | ✓ | ✕ | ✕ |
@@ 4 Review methodology
4.1 Objectives and Research Questions
This research aims to investigate the existing integration of metaheuristics and ML algorithms to detect intrusions in the IoT ecosystem. Besides, uncovering the hidden correlations among the top-performing optimization techniques and ML models, considering specific datasets is another goal of similar importance. To achieve these objectives, firstly, it requires understanding the necessity and exploring different applications of the metaheuristics-assisted ML architectures for developing IoT-IDSs. Secondly, an investigation is needed on the most popular datasets and evaluation metrics utilized to measure these detection systems. The next plan is to discover which optimization techniques and ML models come up with excellent performances, considering the tested datasets. After that, we plan to analyze and sort out the challenges and issues that arise because of the collaboration of these different conceptual techniques. Finally, the unexplored areas of metaheuristics and machine learning techniques need to be studied to facilitate future research. Table 2 outlines the specific research questions identified to achieve the objectives mentioned above.
@@ 4.1 Objectives and Research Questions
4.2 Search strategy
In this review, we search relevant works in established and well-known online sources, such as IEEE Xplore Digital Library, ACM Digital Library, Elsevier, Springer, Nature, Wiley, Taylor & Francis, MDPI, and World Scientific. Initially, 765 articles are selected in total. For searching, specific keywords related to IoT-IDS which incorporate metaheuristics and ML have been used (See Table 3). Since we aim to cover the most recent techniques, we search and select only the relevant, effective, and scholarly papers that were published between 2020 and 2024. Importantly, this focused and extensive investigation differentiates our work from previous SLRs or surveys in the same field. Next, the papers that are deemed irrelevant and found to be duplicates are removed. Particularly, several research has been conducted on the metaheuristics and ML-driven detection systems that incorporate blockchain, fog computing, and supercomputing technologies, which are not relevant to this review. Moreover, several works focus on specific attack detection, for example, botnet, ransomware, malware, denial of service (DoS), etc. These articles are also removed from our dataset. Furthermore, intrusion detection systems that are not IoT-specific, are also eliminated since our concentration is solely on the IoT-IDSs. All of these keywords for exclusion are listed in Table 4. Apart from this, the related surveys are not included in the dataset. Additionally, the papers published in foreign languages are excluded from this study. Finally, 111 articles are chosen for this literature review.
@@ 4.3 Data analysis
4.4 Investigation of Journal Papers
@@ 5 Results of the review
5.1 RQ1: What are the need and existing applications of metaheuristics optimization in developing ML and DL-based IDS, especially for IoT?
Though machine learning algorithms offer high accuracy, enhanced security, and better performance, they require excessive computational power like high-performance GPUs (Graphics Processing Units), and a large volume of storage for generating, executing, and managing large datasets during training and testing phases. Undoubtedly, deep learning algorithms can seamlessly handle enormous datasets and offer fast processing; but, these require significant time to achieve minor precision improvements. Moreover, parameter-tuning is another unavoidable critical issue since adjusting the number of layers with the expected accuracy is entirely correlated [bharati2022machine]. In this regard, metaheuristics can be utilized to obtain near-optimal solutions in a shorter time than the exact algorithms. Specifically, a metaheuristic algorithm is an iterative generation process that guides a subordinate heuristic to explore and exploit the search space efficiently [desale2015heuristic]. Besides, these algorithms terminate when specific conditions (e.g., the number of iterations, elapsed time, etc.) are satisfied. As a result, there is no possibility of running the algorithm for a long period; even the likelihood of being stuck in the local minima is negligible.
@@
5.1.1 Population-based Metaheuristics and ML
@@
Swarm-based Metaheuristics. The well-known particle swarm optimization (PSO) technique is widely used in IoT-IDSs. Particularly, Elmasry et al. [17_elmasry2020evolving] leverage a double PSO to optimize features and hyperparameters. The efficiency of this metaheuristic is examined by discriminative and generative ML methods, especially CNN, LSTM, and DBN. Among these, DBN excels over others by 2% to 6% while testing on the NSL-KDD and CICIDS2017 datasets. In another paper, Saheed et al. [74_saheed2023novel] integrate an AE with PSO to optimally select features from the BoT-IoT and UNSW-NB15 datasets. After that, they also modify the inertia weight of PSO to optimize the hyperparameters of DNN, resulting in the efficient classification of attacks (an accuracy of 97.61% and 94.62% in BoT-IoT and UNSW-NB15, respectively). However, the combination of genetic self-adjusted PSO (GSAPSO) and EGB and KNN classifiers does not result in considerable performances [111_dakic2024intrusion].
@@
Apart from these, bird swarm optimization (BSA), salp swarm algorithm (SSA), and golden jackal optimization algorithm (GJOA) are found to be reliable for detecting intrusions in IoT environments. In particular, a DBN-driven IoT-IDS is presented in [90_biju2024evaluated], which offers 98.96% accuracy, 99.4% precision, and 98.87% recall on the NSL-KDD dataset when parameters are tuned by evaluated BSA (EBSA). Aljehane et al. [94_aljehane2024golden] leverage GJOA to find the most significant features from the CICIDS-2017 dataset and SSA to optimize hyperparameters of the attention-driven bi-directional LSTM (A-BiLSTM). This dual usage of population-based optimizers secures 99.69% accuracy, 98.92% f1-score, and 98.74% MCC. Another variant of SSA, referred to as chaotic salp swarm optimization (CSSA) performs well with the incorporation of LightGBM, having an accuracy of 98.35%99.38% on the MC-IoT, MQTT-IoT-IDS2020, and MQTTset datasets [57_prajisha2022efficient]. However, adopting BSA and social group optimization algorithm (SCOA) for selecting features and optimizing parameters in kernel extreme machine learning model (KELM) fails to deliver decent results across all metrics (99.45% accuracy, 80.26% precision, 82.67% recall, and 80.95% f1-score).
@@
Interestingly, some detection systems demonstrate different performances in various datasets. For example, IDS with artificial bee colony (ABC) and extreme machine learning classifier performs well on CICIDS-2017; whereas significantly fails on UNSW-NB15 (accuracy 98.71% vs. 71.54% ) [34_stankovic2022feature]. Similar results appear in swarm-inspired sand cat swarm optimizer (SCSO) and ELM-based system [50_jovanovic2022feature]. However, integration of glow-swarm optimization (GSO) with PCA [22_anusha2022intrusion], chimp chicken swarm-based optimization (CCSO) with deep LSTM [101_deore2022hybrid], and improved ACO with ensemble classifier (DDT, ANFIS and MDSVM) [39_vanitha2023improved] does not result in considerable performance.
@@
Nature-inspired Metaheuristics. These optimization techniques are extensively integrated with machine learning-based IoT-IDSs, especially Grey-Wolf Optimization (GWO) [56_xu2023iot, 89_elsedimy2024novel, 91_saheed2024voting, 98_moizuddin2022bio, 2_rm2020effective, 61_davahli2020lightweight], firefly optimization algorithm (FOA) [8_savanovic2023intrusion, 60_gangula2023intrusion, 87_harbi2024bio], Capuchin Search Algorithm (CSA) [64_om2022effective, 73_abd2023intrusion], and Whale Optimization Algorithm (WOA) [72_ramana2022wogru, 75_jothi2023wils, 79_mohy2023whale, 80_bajpai2024hybrid]. Regarding the GWO-based detection systems, [56_xu2023iot, 98_moizuddin2022bio, 2_rm2020effective, 61_davahli2020lightweight] utilize GWO for feature selection; whereas [89_elsedimy2024novel, 91_saheed2024voting] use for optimizing hyperparameters of quantum-based SVM classifier and EL method (comprised of DT, RF, KNN, and MLP), respectively. The former systems identify intrusions relying on XGBoost, elastic regularization-assisted contractive autoencoder (CAE), deep neural network (DNN), and SVM, respectively. Notably, all these IDSs come up with remarkable efficiency (e.g., 99%100% accuracy). However, different datasets, preprocessing and other related techniques are employed within these IoT-IDSs. In [60_gangula2023intrusion, 87_harbi2024bio], FOA is integrated for selecting near-optimal features, and detection is conducted by classifiers, especially ensemble, and DT, respectively. Though [60_gangula2023intrusion] shows notable detection ability, [87_harbi2024bio] severely under-performs. In contrast, Savanovic et al. [8_savanovic2023intrusion] improve the original FOA for tuning hyperparameters in classification techniques like KNN, and XGBoost. As a consequence, their IoT-IDS provides an accuracy of 99.98% and 99.6997%, respectively on UNSW-NB15 and IoT healthcare datasets. Turing to CSA-based detection systems, Kumar et al. [64_om2022effective] optimize parameters of a capsule autoencoder (HKCAE) and Elaziz et al. [73_abd2023intrusion] select near-optimal features for CNN model using this metaheuristic. Importantly, in this case, parameter tuning turns to be more effective than feature selection since the former IDS offers outstanding accuracy on BoT-IoT and UNSW-NB15 datasets (99.9% and 99.7%); whereas the later system behaves inconsistently across various datasets (considerable for BoT-IoT, KDDCup-99, and CICIDS-2017; severe for NSL-KDD). Among the four WOA-based IDSs, [72_ramana2022wogru] is the best-performing system with approximately 99.8% accuracy, precision, recall, f1-score, and specificity. The authors leverage this optimization technique to tune the hyperparameters of a gated recurrent unit (GRU). However, optimizing parameters of an LSTM using WOA also turns out to be effective (99.1%99.5% accuracy) [75_jothi2023wils].
@@
Considering other IOT-IDSs of this category, most of the nature-inspired metaheuristics are leveraged for feature selection. Further examining of these systems reveal that various types of optimizers, such as Moth–Flame Optimization (MFO) [66_gadekallu2023moth], crow search algorithm (CrSA) [46_jayalatchumy2024improved], chaotic vortex search (CVS) [51_geetha2024cvs], decisive red fox optimization (DRFO) [88_rabie2024novel], multi-objective prairie dog optimization (PDO [92_sharmamulti], reptile search algorithm (RSA) [16_dahou2022intrusion], binary multi-objective Capuchin search algorithm (BMECapSA) [42_asgharzadeh2023anomaly], Aquila optimizer (AQUO) [49_fatani2021advanced], BA [95_bella2024intrusion], and Mayfly Optimization Algorithm (MOA) [96_vadakkethil2024mayfly] are prominent. Among them, [66_gadekallu2023moth, 46_jayalatchumy2024improved, 51_geetha2024cvs, 88_rabie2024novel, 92_sharmamulti, 42_asgharzadeh2023anomaly] demonstrate higher efficiency while experimenting on diverse datasets. However, no common trend is found in terms of machine learning-based classifiers.
@@
Other than these systems, a few IDSs utilize nature-centric optimization techniques for tuning parameters [93_cherian2024iot, 53_qaddoura2024evolving, 82_kethineni2024intrusion] or hyperparameters [85_vijayan2024original, 114_alamro2023modeling] of the detection models; sometimes for hybrid applications [78_gopalakrishnan2022new, 100_pingale2023remora, 106_escorcia2023sea]. Although in discrete cases, these systems show considerable results, their performances are not generally satisfactory.
@@
Evolutionary Algorithms. Only a small number of IoT-IDSs have utilized evolutionary algorithms (EAs) and they do not significantly surpass other detection systems. The best-performing IDS of this type is proposed by Latif et al. [112_latif2024dtl], where hyperparameters of a CNN-based ensemble classifier are optimized by a genetic algorithm (GA). Ultimately, this system achieves 100% accuracy, precision, recall, f1-score, and Cohen’s kappa score in the Edge_IIoTset dataset. Second best-performing IDSs presented in [12_dey2023hybrid], where feature selection is accomplished by a non-dominated sorting genetic algorithm (NSGA) and classification is done using a support vector machine (SVM). Importantly, the system achieves a remarkable accuracy of 99.48% on the TON_IoT dataset. Gupta et al. [104_10317882] integrate an evolutionary algorithm intelligent water drop (IWD) and a nature-inspired biogeography-based optimization (BBO) technique with a feed-forward neural network (FNN). The IDS detects attacks more correctly when tested on CICIDS-2017 compared to the IoTID20 dataset (accuracy 98.2339% vs. 96.7414% and f1-score 99.0865% vs. 95.4901%). However, integration of an assimilated artificial fish swarm optimization (AAFSO) with genetic algorithm (GA)-tuned faster recurrent CNN (FRCNN) does not perform well across diverse datasets [65_anushiya2023new].

### v4

@@ III Related Work
The systematic literature review (SLR) in [82] covers only a few population-based metaheuristics for intrusion detection in the IoT environment. Additionally, physical law-based, human-inspired, and hybrid optimization techniques are not discussed. Moreover, the authors do not concentrate solely on IoT. Rather, wireless, public networks, computer networks, Hadoop and MapReduce, and edge networks are also explored significantly. Furthermore, though they analyze the datasets used in developing the IDSs, no correlation is discovered among the metaheuristics, mostly used datasets and ML methods. Saadouni et al. [143] presents an SLR for IoT-IDS based on bio-inspired and ML-driven techniques. The authors vividly discuss the integration of ML methods with optimization algorithms. However, one of the most notable limitations of this work is the investigation of only 25 papers, whereas there are several well-established metaheuristics-assisted IDSs dedicated to IoT. Importantly, although the authors claim to study only IoT-based papers, we find that most of the articles are not focused on the IoT environment. In [141], the authors study population-based optimizations, specifically swarm intelligence devised for detecting intrusions in IoT. Besides, they analyze the datasets used and the performances of the existing systems. The main drawback of the SLR is the lack of covering all categories of metaheuristics-driven IDSs. Moreover, only basic ML algorithms are discussed in the SLR; whereas many crucial deep learning-based IDSs are sorted out in our study. Sharma et al. [157] aims to explore IoT-IDSs that rely on only multi-objective metaheuristics algorithms. Apart from this, they analyze different machine-learning models and popular datasets. However, the relation between these diverse techniques is still missing. Moreover, the study is not systematic. Heidari et al. [80] introduces an SLR containing rigid comparison and exploration of different IDSs in the IoT environment. The first and foremost limitation is the missing metaheuristics and ML-based systems, which we aim to cover in our study. Verma et al. [169] propose a survey on ML-driven IDSs for IoT applications. Regrettably, the study does not include any optimization algorithms, rather it intends to analyze the machine learning classifiers commonly utilized in intrusion detection. Hajiheida et al. [76] also do not study the metaheuristics algorithms that are extensively employed in the IoT-IDS. Importantly, rather than focusing on the metaheuristics or ML-based systems, they categorize and discuss different systems, such as SIDS, AIDS, SpIDS, and HyIDS.
@@
The existing reviews either explore nature-inspired or ML-based IoT-IDS systems. Though there is only one survey [157] that discusses the integration of metaheuristics and machine learning techniques for the IoT environment, the coverage and analysis are too inadequate regarding the volume, significance, and diversity of the related IDSs in literature. Most importantly, no reviews analyze the correlation between the optimization algorithms and machine learning models while experimenting on a specific dataset. Apart from these, the existing works do not categorize the applications of different metaheuristics regarding intrusion detection in the IoT environment.
@@
To address all these issues, a systematic literature review is presented, which extensively explores the existing metaheuristics and ML-integrated IDSs, specific to the IoT environment. Additionally, we analyze these systems based on different applications of metaheuristics algorithms, such as optimal feature selection, parameter tuning, hyperparameter tuning, etc. Moreover, the discovery of the connections among these algorithms, their outstanding performances, and the used datasets significantly distinguish this review from others. Furthermore, a new large-scale visualized taxonomy is demonstrated to provide researchers with an overview of the existing IoT-IDSs. Table I provides an in-depth comparison with the state-of-the-art reviews, highlighting the contributions of our work over others.
@@
| [82] | 2024 | ✓ | 145 | ✕ | ✕ | ✓ | ✕ | ✕ | ✕ | ✕ | ✕ | ✕ | ✓ | ✕ | ✕ |
| [143] | 2024 | ✓ | 25 | ✕ | ✓ | ✓ | ✕ | ✕ | ✓ | ✓ | ✕ | ✕ | ✓ | ✕ | ✕ |
| [141] | 2024 | ✓ | 101 | ✕ | ✕ | ✓ | ✕ | ✕ | ✓ | ✕ | ✕ | ✓ | ✓ | ✕ | ✕ |
| [157] | 2024 | ✕ | 37 | ✓ | ✓ | ✓ | ✕ | ✓ | ✓ | ✓ | ✕ | ✕ | ✕ | ✕ | ✕ |
| [80] | 2023 | ✓ | 24 | ✓ | ✕ | ✕ | ✕ | ✕ | ✕ | ✕ | ✕ | ✕ | ✓ | ✕ | ✕ |
| [169] | 2020 | ✕ | 25* | ✓ | ✕ | ✕ | ✕ | ✕ | ✓ | ✕ | ✕ | ✕ | ✓ | ✕ | ✕ |
| [76] | 2019 | ✓ | 43 | ✓ | ✕ | ✕ | ✕ | ✕ | ✓ | ✓ | ✕ | ✕ | ✓ | ✕ | ✕ |
@@ IV Review methodology
IV-A Objectives and Research Questions
This research aims to investigate the existing integration of metaheuristics and ML algorithms to detect intrusions in the IoT ecosystem. Besides, uncovering the hidden correlations among the top-performing optimization techniques and ML models, considering specific datasets is another goal of similar importance. To achieve these objectives, firstly, it requires understanding the necessity and exploring different applications of the metaheuristics-assisted ML architectures for developing IoT-IDSs. Secondly, an investigation is needed on the most popular datasets and evaluation metrics utilized to measure these detection systems. The next plan is to discover which optimization techniques and ML models come up with excellent performances, considering the tested datasets. After that, we plan to analyze and sort out the challenges and issues that arise because of the collaboration of these different conceptual techniques. Finally, the unexplored areas of metaheuristics and machine learning techniques need to be studied to facilitate future research. Table II outlines the specific research questions identified to achieve the objectives mentioned above.
@@ IV-A Objectives and Research Questions
IV-B Search strategy
In this review, we search relevant works in established and well-known online sources, such as IEEE Xplore Digital Library, ACM Digital Library, Elsevier, Springer, Nature, Wiley, Taylor & Francis, MDPI, and World Scientific. Initially, 765 articles are selected in total. For searching, specific keywords related to IoT-IDS which incorporate metaheuristics and ML have been used (See Table III). Since we aim to cover the most recent techniques, we search and select only the relevant, effective, and scholarly papers that were published between 2020 and 2024. Importantly, this focused and extensive investigation differentiates our work from previous SLRs or surveys in the same field. Next, the papers that are deemed irrelevant and found to be duplicates are removed. Particularly, several research has been conducted on the metaheuristics and ML-driven detection systems that incorporate blockchain, fog computing, and supercomputing technologies, which are not relevant to this review. Moreover, several works focus on specific attack detection, for example, botnet, ransomware, malware, denial of service (DoS), etc. These articles are also removed from our dataset. Furthermore, intrusion detection systems that are not IoT-specific, are also eliminated since our concentration is solely on the IoT-IDSs. All of these keywords for exclusion are listed in Table IV. Apart from this, the related surveys are not included in the dataset. Additionally, the papers published in foreign languages are excluded from this study. Finally, 111 articles are chosen for this literature review.
@@ IV-C Data analysis
IV-D Investigation of Journal Papers
@@ V Results of the review
V-A RQ1: What are the need and existing applications of metaheuristics optimization in developing ML and DL-based IDS, especially for IoT?
Though machine learning algorithms offer high accuracy, enhanced security, and better performance, they require excessive computational power like high-performance GPUs (Graphics Processing Units), and a large volume of storage for generating, executing, and managing large datasets during training and testing phases. Undoubtedly, deep learning algorithms can seamlessly handle enormous datasets and offer fast processing; but, these require significant time to achieve minor precision improvements. Moreover, parameter-tuning is another unavoidable critical issue since adjusting the number of layers with the expected accuracy is entirely correlated [33]. In this regard, metaheuristics can be utilized to obtain near-optimal solutions in a shorter time than the exact algorithms. Specifically, a metaheuristic algorithm is an iterative generation process that guides a subordinate heuristic to explore and exploit the search space efficiently [44]. Besides, these algorithms terminate when specific conditions (e.g., the number of iterations, elapsed time, etc.) are satisfied. As a result, there is no possibility of running the algorithm for a long period; even the likelihood of being stuck in the local minima is negligible.
@@
V-A1 Population-based Metaheuristics and ML
@@
Swarm-based Metaheuristics. The well-known particle swarm optimization (PSO) technique is widely used in IoT-IDSs. Particularly, Elmasry et al. [54] leverage a double PSO to optimize features and hyperparameters. The efficiency of this metaheuristic is examined by discriminative and generative ML methods, especially CNN, LSTM, and DBN. Among these, DBN excels over others by 2% to 6% while testing on the NSL-KDD and CICIDS2017 datasets. In another paper, Saheed et al. [146] integrate an AE with PSO to optimally select features from the BoT-IoT and UNSW-NB15 datasets. After that, they also modify the inertia weight of PSO to optimize the hyperparameters of DNN, resulting in the efficient classification of attacks (an accuracy of 97.61% and 94.62% in BoT-IoT and UNSW-NB15, respectively). However, the combination of genetic self-adjusted PSO (GSAPSO) and EGB and KNN classifiers does not result in considerable performances [40].
@@
Apart from these, bird swarm optimization (BSA), salp swarm algorithm (SSA), and golden jackal optimization algorithm (GJOA) are found to be reliable for detecting intrusions in IoT environments. In particular, a DBN-driven IoT-IDS is presented in [34], which offers 98.96% accuracy, 99.4% precision, and 98.87% recall on the NSL-KDD dataset when parameters are tuned by evaluated BSA (EBSA). Aljehane et al. [12] leverage GJOA to find the most significant features from the CICIDS-2017 dataset and SSA to optimize hyperparameters of the attention-driven bi-directional LSTM (A-BiLSTM). This dual usage of population-based optimizers secures 99.69% accuracy, 98.92% f1-score, and 98.74% MCC. Another variant of SSA, referred to as chaotic salp swarm optimization (CSSA) performs well with the incorporation of LightGBM, having an accuracy of 98.35%99.38% on the MC-IoT, MQTT-IoT-IDS2020, and MQTTset datasets [134]. However, adopting BSA and social group optimization algorithm (SCOA) for selecting features and optimizing parameters in kernel extreme machine learning model (KELM) fails to deliver decent results across all metrics (99.45% accuracy, 80.26% precision, 82.67% recall, and 80.95% f1-score).
@@
Interestingly, some detection systems demonstrate different performances in various datasets. For example, IDS with artificial bee colony (ABC) and extreme machine learning classifier performs well on CICIDS-2017; whereas significantly fails on UNSW-NB15 (accuracy 98.71% vs. 71.54% ) [161]. Similar results appear in swarm-inspired sand cat swarm optimizer (SCSO) and ELM-based system [89]. However, integration of glow-swarm optimization (GSO) with PCA [21], chimp chicken swarm-based optimization (CCSO) with deep LSTM [43], and improved ACO with ensemble classifier (DDT, ANFIS and MDSVM) [167] does not result in considerable performance.
@@
Nature-inspired Metaheuristics. These optimization techniques are extensively integrated with machine learning-based IoT-IDSs, especially Grey-Wolf Optimization (GWO) [173, 55, 145, 124, 142, 41], firefly optimization algorithm (FOA) [152, 67, 78], Capuchin Search Algorithm (CSA) [128, 1], and Whale Optimization Algorithm (WOA) [140, 88, 123, 27]. Regarding the GWO-based detection systems, [173, 124, 142, 41] utilize GWO for feature selection; whereas [55, 145] use for optimizing hyperparameters of quantum-based SVM classifier and EL method (comprised of DT, RF, KNN, and MLP), respectively. The former systems identify intrusions relying on XGBoost, elastic regularization-assisted contractive autoencoder (CAE), deep neural network (DNN), and SVM, respectively. Notably, all these IDSs come up with remarkable efficiency (e.g., 99%100% accuracy). However, different datasets, preprocessing and other related techniques are employed within these IoT-IDSs. In [67, 78], FOA is integrated for selecting near-optimal features, and detection is conducted by classifiers, especially ensemble, and DT, respectively. Though [67] shows notable detection ability, [78] severely under-performs. In contrast, Savanovic et al. [152] improve the original FOA for tuning hyperparameters in classification techniques like KNN, and XGBoost. As a consequence, their IoT-IDS provides an accuracy of 99.98% and 99.6997%, respectively on UNSW-NB15 and IoT healthcare datasets. Turing to CSA-based detection systems, Kumar et al. [128] optimize parameters of a capsule autoencoder (HKCAE) and Elaziz et al. [1] select near-optimal features for CNN model using this metaheuristic. Importantly, in this case, parameter tuning turns to be more effective than feature selection since the former IDS offers outstanding accuracy on BoT-IoT and UNSW-NB15 datasets (99.9% and 99.7%); whereas the later system behaves inconsistently across various datasets (considerable for BoT-IoT, KDDCup-99, and CICIDS-2017; severe for NSL-KDD). Among the four WOA-based IDSs, [140] is the best-performing system with approximately 99.8% accuracy, precision, recall, f1-score, and specificity. The authors leverage this optimization technique to tune the hyperparameters of a gated recurrent unit (GRU). However, optimizing parameters of an LSTM using WOA also turns out to be effective (99.1%99.5% accuracy) [88].
@@
Considering other IOT-IDSs of this category, most of the nature-inspired metaheuristics are leveraged for feature selection. Further examining of these systems reveal that various types of optimizers, such as Moth–Flame Optimization (MFO) [66], crow search algorithm (CrSA) [85], chaotic vortex search (CVS) [68], decisive red fox optimization (DRFO) [137], multi-objective prairie dog optimization (PDO [156], reptile search algorithm (RSA) [39], binary multi-objective Capuchin search algorithm (BMECapSA) [24], Aquila optimizer (AQUO) [60], BA [32], and Mayfly Optimization Algorithm (MOA) [165] are prominent. Among them, [66, 85, 68, 137, 156, 24] demonstrate higher efficiency while experimenting on diverse datasets. However, no common trend is found in terms of machine learning-based classifiers.
@@
Other than these systems, a few IDSs utilize nature-centric optimization techniques for tuning parameters [37, 136, 97] or hyperparameters [170, 8] of the detection models; sometimes for hybrid applications [71, 133, 56]. Although in discrete cases, these systems show considerable results, their performances are not generally satisfactory.
@@
Evolutionary Algorithms. Only a small number of IoT-IDSs have utilized evolutionary algorithms (EAs) and they do not significantly surpass other detection systems. The best-performing IDS of this type is proposed by Latif et al. [106], where hyperparameters of a CNN-based ensemble classifier are optimized by a genetic algorithm (GA). Ultimately, this system achieves 100% accuracy, precision, recall, f1-score, and Cohen’s kappa score in the Edge_IIoTset dataset. Second best-performing IDSs presented in [46], where feature selection is accomplished by a non-dominated sorting genetic algorithm (NSGA) and classification is done using a support vector machine (SVM). Importantly, the system achieves a remarkable accuracy of 99.48% on the TON_IoT dataset. Gupta et al. [73] integrate an evolutionary algorithm intelligent water drop (IWD) and a nature-inspired biogeography-based optimization (BBO) technique with a feed-forward neural network (FNN). The IDS detects attacks more correctly when tested on CICIDS-2017 compared to the IoTID20 dataset (accuracy 98.2339% vs. 96.7414% and f1-score 99.0865% vs. 95.4901%). However, integration of an assimilated artificial fish swarm optimization (AAFSO) with genetic algorithm (GA)-tuned faster recurrent CNN (FRCNN) does not perform well across diverse datasets [23].